# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v2

@@ Abstract
FedProx is a simple yet effective federated learning method that enables model personalization via regularization. Despite remarkable success in practice, a rigorous analysis of how such a regularization provably improves the statistical accuracy of each client’s local model hasn’t been fully established. Setting the regularization strength heuristically presents a risk, as an inappropriate choice may even degrade accuracy. This work fills in the gap by analyzing the effect of regularization on statistical accuracy, thereby providing a theoretical guideline for setting the regularization strength for achieving personalization. We prove that by adaptively choosing the regularization strength under different statistical heterogeneity, FedProx can consistently outperform pure local training and achieve a nearly minimax-optimal statistical rate. In addition, to shed light on resource allocation, we design an algorithm, provably showing that stronger personalization reduces communication complexity without increasing the computation cost overhead. Finally, our theory is validated on both synthetic and real-world datasets and its generalizability is verified in a non-convex setting.
Keywords: Personalized Federated Learning, Statistical Complexity, Communication Complexity, Minimax Optimality
@@ 1 Introduction
$\bullet$ Nearly Optimal Statistical Accuracy. We analyze the statistical accuracy of the objective in (1), taking statistical heterogeneity into account. For each of the models, we establish a state-of-the-art convergence rate $\mathcal{O}(1/(mn))+\mathcal{O}(1/n\wedge R/n^{1/2})$, where $R$ characterizes the statistical heterogeneity and $n$ is the sample size on each client. The first term $\mathcal{O}(1/(mn))$ represents the sample complexity of training a single global model under homogeneity, which is attained by (1) when $R$ is close to zero, highlighting the benefit of FL. For the second term, the first component $\mathcal{O}(1/n)$ is the sample complexity of purely local training and the second component $\mathcal{O}(R/n^{1/2})$ demonstrates the flexibility of FedProx under different statistical heterogeneity, implying the statistical accuracy of FedProx never gets worse than purely local training’s. The fine-grained analysis demonstrates that with an adaptive choice of personalization degree characterized by $\lambda$, FedProx achieves a nearly minimax-optimal statistical accuracy across different levels of statistical heterogeneity.
$\bullet$ Provably Enhanced Communication Efficiency. As the objective (1) could be reformulated as a bi-level problem, we propose an algorithm for (1) with an explicit computation cost assessment. We establish the communication complexity as $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log\frac{1}{\varepsilon})$ and the local computation complexity in each communication round as $\mathcal{\widetilde{O}}({\frac{\lambda+L}{\lambda+\mu}})$. With higher personalization degree, communication complexity would decrease and the computation complexity per round would increase, demonstrating that there is a trade-off between computation and communication costs. Our analysis concludes that model personalization can reduce communication overhead without incurring additional computation expense.
@@ 4 When will Personalization Improve Statistical Accuracy
Intuitively, a larger $R$ indicates a greater divergence between the clients’ data distributions, thus stronger statistical heterogeneity, and vice versa. Therefore, in the FedProx objective, the choice of $\lambda$ should be adjusted according to $R$. Next, we introduce two standard assumptions.
@@
Each loss function $L_{i}$ is $\mu$-strongly convex, i.e., $L_{i}({\bm{w}},S_{i})\geq L_{i}({\bm{w}}^{\prime},S_{i})+\langle\nabla L_{i}({\bm{w}}^{\prime},S_{i}),{\bm{w}}-{\bm{w}}^{\prime}\rangle+\frac{\mu}{2}\|{\bm{w}}-{\bm{w}}^{\prime}\|^{2},\quad\forall{\bm{w}},{\bm{w}}^{\prime}\in\mathbb{R}^{d},i\in[m].$
Assumption 3 (Bound Gradient Variance at the Optimum).
@@
Here, the strongly convex assumption is used to establish the estimation error and is widely adopted in the statistical analysis of regularization-based personalized federated learning (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and statistical analysis in federated learning (Chen et al., 2023b; Cheng et al., 2023). Assumption 3 is used to quantify the observation noise, and is standard in the literature (Li et al., 2020b; Chen et al., 2023b).
Next, we establish a statistical rate of the optimal solution to the Fedprox problem (1) with a fine-grained analysis. To facilitate the analysis, we first define some notations. Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models. A check on the optimality condition implies that the optimal solution for the global model is $\widetilde{\bm{w}}^{(g)}=\sum^{m}_{i=1}p_{i}\widetilde{\bm{w}}^{(i)}$. In this paper, we use global statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and local statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as the measure of statstical accuracy. Here the expectation is taken with respect to all data. Note that some existing literature adopts alternative metrics, such as average excess risk and individual excess risk (Chen et al., 2023b) to measure the statistical error. They are equivalent to our metrics under the assumptions used in this paper. We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error. See A.2 for the proof.
@@
Suppose Assumptions 1 and 2 hold. We have
$\underbrace{\sum_{i\in[m],j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\mu\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\rangle\right)}_{\text{model estimation error}}\leq\frac{\lambda}{2}\underbrace{(R^{2}-\sum_{i\in[m]}p_{i}\|\widetilde{\bm{w}}^{(i)}-\widetilde{\bm{w}}^{(g)}\|^{2})}_{\text{heterogeneity estimation error}}.$ (3)
Lemma 1 establishes a relationship between the local model estimation error and the statistical heterogeneity. On the right-hand side of (3), the first term $R$, which is larger than $\|{{\bm{w}}_{\star}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$, measures the statistical heterogeneity; the second term, $\sum_{i=1}^{m}p_{i}\|\widetilde{\bm{w}}^{(i)}-\widetilde{\bm{w}}^{(g)}\|^{2}$, can be understood as the estimator of such heterogeneity. Lemma 1 implies that to achieve a tight bound on the estimation error, we should choose a $\lambda$ that “minimizes” the right-hand side of (3). When $R$ is large, one straightforward choice is $\lambda=\mathcal{O}(1/(nR^{2}))$, which will be shown to yield an error bound of order $\mathcal{O}(1/n)$, matching the performance of purely local training and is independent of $R$. When $R$ is small, $\lambda$ should be carefully chosen to obtain the full sample complexity $\mathcal{O}(1/(mn))$. In the following theorem, we provide such a choice and establish a state-of-the-art statistical rate.
@@
Suppose Assumptions 1, 2 and 3 hold and $n_{i}=n,\forall i\in[m]$. Consider the FedProx objective in (1) with $p_{i}=\frac{1}{m},\forall i\in[m]$. If $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$, the global model $\widetilde{\bm{w}}^{(g)}$ and local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained from solving the objective (1) satisfy
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\right\|^{2}\leq\frac{C_{1}}{N}+C_{2}\left(\frac{1}{n}\wedge\frac{R}{\sqrt{n}}\right),$ (4)
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq\frac{C_{3}}{N}+C_{4}(\frac{1}{n}\wedge\frac{R}{\sqrt{n}})+16(\frac{\rho^{2}}{\mu^{2}n}\wedge R^{2})\quad\text{ for all }i\in[m],$ (5)
@@
$\begin{split}C_{1}&=\frac{\rho^{2}}{\mu^{2}}=\frac{C_{3}}{8},C_{2}=\big{(}\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\big{)}^{2}\vee\frac{2\rho}{\mu},C_{4}=\big{(}\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\big{)}^{2}\wedge(8C_{2}+\frac{2\rho}{\mu}).\end{split}$ (6)
@@
See Appendix A.2 for (4) and Appendix A.3 for (5). ∎
Theorem 1 shows that with an appropriate choice of $\lambda$, both the global and local models achieve a fast statistical rate across a wide range of $R$. To interpret the result, first we analyze the error bound provided in (4) and (5). The two error bounds have the same order and can be decomposed into two key components. The first term, of order $\mathcal{O}(N^{-1})$, captures the benefit of collaborative learning, as $N=mn$ is the total sample size across $m$ clients. This term reflects the sample efficiency gained through collaborative learning. The second term, of order $\mathcal{O}(n^{-1}\wedge n^{-1/2}R)$, measures the cost of statistical heterogeneity. As $R$ increases, the convergence rate gradually degrades, transitioning from $\mathcal{O}(N^{-1})$ to $\mathcal{O}(n^{-1})$, showing that the error bound of FedProx interpolates between the two extremes: LocalTrain and GlobalTrain.
Next, to better understand how such an interpolation is achieved through an adaptive personalization degree, we may examine the two extreme cases: as the statistical heterogeneity $R\to\infty$, and $\lambda\to 0$, leading to a high degree of personalization. The statistical error bound converges to $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, and $\lambda\to\infty$, leading to a low degree of personalization. The statistical error bound converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain. As $R\in(0,\infty)$, $\lambda$ is adjusted correspondingly, making FedProx always perform no worse than LocalTrain while benefiting from global training when clients share similarities.
@@
(Chen et al., 2023b) A1,2,3, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
Ours A1,2,3 One-stage $\frac{1}{n}$ $\frac{1}{\sqrt{n}}R$ $\frac{1}{mn}$
Lower Bound - - $\frac{1}{n}$ $R^{2}\vee\frac{1}{mn}$ $\frac{1}{mn}$
We now compare our theoretical results with the result up to date (Chen et al., 2023b), as summarized in Table 1. Our contribution can be summarized into three aspects:
@@
$\bullet$ We establish a state-of-the-art nearly minimax statistical rate. As shown in Table 1, in a heterogeneous case when $R\geq n^{-1/2}$, Chen et al. (2023b) establishes a rate that is at least of the order $\mathcal{O}\left(n^{-1}R^{2}\right)$ as $D\geq R$. This bound becomes increasingly loose with larger $R$. In contrast, our bound is $\mathcal{O}\left(n^{-1}\right)$, which shows that properly-tuned FedProx is always no worse than LocalTrain, independent of $R$. In a homogeneous case when $R\leq m^{-1}n^{-1/2}$, they can only establish a rate slower than $\mathcal{O}\left(1/(\sqrt{m}n)\right)$, while our result achieves a rate of $\mathcal{O}\left(1/(mn)\right)$, leveraging all $mn$ samples and matching the rate of GlobalTrain on the IID data. Compared to the established lower bound (Chen et al., 2018), our result is nearly minimax optimal. Specifically, when $R>n^{-1/2}$ and $R\leq m^{-1}n^{-1/2}$, our result matches the lower bound. There is a small gap exists in the case when $m^{-1}n^{-1/2}<R\leq n^{-1/2}$, where we achieve a rate of $\mathcal{O}(n^{-1/2}R)$ whereas the lower bound is $\mathcal{O}(R^{2})$. Such a gap is due to the cost of estimating the model difference $\{{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\}_{i=1}^{m}$ when $\lambda$ is finite (c.f. Lemma 1). The improvement of this gap is non-trivial, which we defer to future work.

### v3

@@ Abstract
FedProx is a simple yet effective federated learning method that enables model personalization via regularization. Despite remarkable success in practice, a rigorous analysis of how such a regularization provably improves the statistical accuracy of each client’s local model hasn’t been fully established. Setting the regularization strength heuristically presents a risk, as an inappropriate choice may even degrade accuracy. This work fills in the gap by analyzing the effect of regularization on statistical accuracy, thereby providing a theoretical guideline for setting the regularization strength for achieving personalization. We prove that by adaptively choosing the regularization strength under different statistical heterogeneity, FedProx can consistently outperform pure local training and achieve a minimax-optimal statistical rate. In addition, to shed light on resource allocation, we design an algorithm, provably showing that stronger personalization reduces communication complexity without increasing the computation cost overhead. Finally, our theory is validated on both synthetic and real-world datasets and its generalizability is verified in a non-convex setting.
@@ 1 Introduction
$\bullet$ Minimax-optimal Statistical Accuracy. We analyze the statistical accuracy of the objective in (1), taking statistical heterogeneity into account. For each of the models, we establish a state-of-the-art convergence rate $\mathcal{O}(1/(mn))+\mathcal{O}(1/n\wedge R^{2})$, where $R$ characterizes the statistical heterogeneity and $n$ is the sample size on each client. The first term $\mathcal{O}(1/(mn))$ represents the sample complexity of training a single global model under homogeneity, which is attained by (1) when $R$ is close to zero, highlighting the benefit of FL. For the second term, the first component $\mathcal{O}(1/n)$ is the sample complexity of purely local training, and the second component $\mathcal{O}(R^{2})$ demonstrates the flexibility of FedProx under different statistical heterogeneity, implying the statistical accuracy of FedProx never gets worse than purely local training’s. Such a rate matches the existing lower bound for the problem. Therefore, to the best of our knowledge, this is the first work that demonstrates that with a theoretically-informed adaptive choice of personalization degree characterized by $\lambda$, the objective in (1) achieves a minimax-optimal statistical accuracy across different levels of statistical heterogeneity. The theoretical analysis we use for objective (1) would extend to different objectives in regularization-based personalized federated learning (Smith et al., 2017; Li et al., 2020a; Hanzely & Richtárik, 2020; Hanzely et al., 2020; 2023).
$\bullet$ Provably Enhanced Communication Efficiency. As the objective (1) could be reformulated as a bi-level problem, we propose an algorithm for (1) with an explicit computation cost assessment. We establish the communication complexity as $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log\frac{1}{\varepsilon})$ and the local computation complexity in each communication round as $\mathcal{\widetilde{O}}({\frac{\lambda+L}{\lambda+\mu}})$. With higher personalization degree, communication complexity would decrease and the computation complexity per round would increase, demonstrating that there is a trade-off between computation and communication costs. Our analysis concludes that model personalization can reduce communication overhead without incurring additional computation expenses.
@@ 4 When will Personalization Improve Statistical Accuracy
Intuitively, a larger $R$ indicates a greater divergence between the clients’ data distributions, thus stronger statistical heterogeneity, and vice versa. Therefore, in the FedProx objective, the choice of $\lambda$ should be adjusted according to $R$. More discussion about Assumption 1 can be found in Appendix A.1. Next, we introduce three standard assumptions.
Assumption 2 (Smoothness).
@@
The loss function $\ell(\cdot,z)$ is almost surely $\mu$-strongly convex: $\ell(x,z)\geq\ell(y,z)+\langle\nabla\ell(y,z),y-x\rangle+\frac{\mu}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{d}.$
Assumption 4 (Bound Gradient Variance at Optimum).
@@
Here, the strongly convex and smooth assumption is used to establish the estimation error and is widely adopted in the theoretical analysis of regularization-based personalized federated learning (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and theoretical analysis in federated learning (Chen et al., 2023b; Cheng et al., 2023). Assumption 4 is used to quantify the observation noise, and is also standard in the literature (Duan & Wang, 2023; Chen et al., 2023b).
Next, we establish a statistical rate of the optimal solution to the Fedprox problem (1) with a fine-grained analysis. To facilitate the analysis, we first define some notations. Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models $\bm{w}^{(i)}$ in objective (1). A check on the optimality condition implies that the optimal solution for the global model is $\widetilde{\bm{w}}^{(g)}=\sum^{m}_{i=1}p_{i}\widetilde{\bm{w}}^{(i)}$. In this paper, we use global statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and local statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as the measure of statstical accuracy. Here the expectation is taken with respect to all data. Note that some existing literature adopts alternative metrics, such as average excess risk and individual excess risk (Chen et al., 2023b) to measure the statistical error. The convergence rate of these two definitions is closely related up to the smoothness constant and strong convexity constant. Define $\tilde{\bm{w}}_{\text{GT}}$ as the optimal solution for the GlobalTrain problem in (2). We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error. See Appendix A.2 for the proof.
@@
Suppose Assumptions 1, 2 and 3 hold. We have
$\|\widetilde{\bm{w}}^{(g)}-\tilde{\bm{w}}_{\text{GT}}\|^{2}\leq\frac{4\left(1+L/\mu\right)}{\mu(\mu+\lambda)}\sum_{i\in[m]}p_{i}\|\nabla L_{i}(\widetilde{\bm{w}}^{(g)},S_{i})\|^{2},$ (4)
$\sum_{i\in[m],j\in[n_{i}]}\frac{p_{i}}{n_{i}}$ $\left(\mu\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\rangle\right)\leq\frac{\lambda}{2}R^{2}.$ (5)
Lemma 1 establishes a relationship between the model estimation error and the tuning parameter $\lambda$. From the bound (4), we can see that a larger $\lambda$ will push the global model solution $\widetilde{\bm{w}}^{(g)}$ towards the GlobalTrain solution. On the other hand, from the bound (5), we can set a smaller $\lambda\propto 1/nR^{2}$ to obtain a rate $O(1/n)$, independent of $R$. This suggests that $\lambda$ can be chosen adaptively w.r.t. $R$ to achieve the optimal statistical rate. A detailed discussion can be found in Appendix A.3. In the following theorem, we provide such a choice and establish a state-of-the-art statistical rate.
@@
Suppose Assumptions 1, 2, 3 and 4 hold and $n_{i}=n,\forall i\in[m]$. Consider the FedProx objective in (1) with $p_{i}=\frac{1}{m},\forall i\in[m]$. If it takes $\lambda\geq\max\left\{64\kappa^{2}L,\left(2\kappa\vee 5\right)\mu\frac{2L^{2}R^{2}+\rho^{2}/n}{L^{2}R^{2}+\rho^{2}/N}\right\}-\mu$ when $R\leq\frac{1}{\sqrt{n}}$, and $\lambda\leq\frac{\rho^{2}}{n\mu R^{2}}$ when $R>\frac{1}{\sqrt{n}}$, the global model $\widetilde{\bm{w}}^{(g)}$ and local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained from solving the objective (1) satisfy
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\right\|^{2}\leq\frac{C_{1}}{N}+C_{2}\left(\frac{1}{n}\wedge R^{2}\right),$ (6)
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq\frac{C_{3}}{N}+C_{4}(\frac{1}{n}\wedge R^{2})\quad\text{ forall }i\in[m],$ (7)
@@
$C_{1}=32\frac{\rho^{2}}{\mu^{2}},C_{2}=\big{(}\frac{(1+\sqrt{3})\rho}{\mu}\big{)}^{2}\vee\frac{32L^{2}}{\mu^{2}},C_{3}=\frac{192\rho^{2}}{\mu^{2}},C_{4}=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee\frac{198L^{2}}{\mu^{2}}.$
@@
See Appendix A.3. ∎
Theorem 1 shows that with an appropriate choice of $\lambda$, both the global and local models achieve a fast statistical rate across a wide range of $R$. To interpret the result, we first analyze the error bound provided in (6) and (7). The two error bounds have the same order and can be decomposed into two key components. The first term, of order $\mathcal{O}(N^{-1})$, captures the benefit of collaborative learning, as $N=mn$ is the total sample size across $m$ clients. This term reflects the sample efficiency gained through collaborative learning. The second term, of order $\mathcal{O}(n^{-1}\wedge R^{2})$, measures the cost of statistical heterogeneity. As $R$ increases, the convergence rate gradually degrades, transitioning from $\mathcal{O}(N^{-1})$ to $\mathcal{O}(n^{-1})$. Notably, the guarantee is minimax-optimal, as the derived bound matches the problem’s lower bound under the given assumptions (Chen et al., 2023b). Remark 2 in Appendix A.3 discusses about how such an interpolation is achieved through different personalization degrees.
@@
Chen et al. (2023b) A1,2,3,4, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
Ours A1,2,3,4 One-stage $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
Lower Bound - - $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
We now compare our theoretical results with the result up to date (Chen et al., 2023b), as summarized in Table 1. For a fair comparison, it is noticeable that we have transformed their result into our metrics $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, by removing their optimization error and problem-dependent constant. Our contribution can be summarized into three aspects:
@@
$\bullet$ We establish a state-of-the-art minimax statistical rate. As shown in Table 1, the established upper bound matches the lower bound established in Chen et al. (2023b). In a heterogeneous case when $R\geq n^{-1/2}$, Chen et al. (2023b) establishes a rate that is at least of the order $\mathcal{O}\left(n^{-1}R^{2}\right)$ as $D\geq R$. This bound becomes increasingly loose with larger $R$. In contrast, our bound is $\mathcal{O}\left(n^{-1}\right)$, which shows that properly-tuned FedProx is always no worse than LocalTrain, independent of $R$. In a homogeneous case when $R\leq m^{-1}n^{-1/2}$, they can only establish a rate slower than $\mathcal{O}\left(1/(\sqrt{m}n)\right)$, while our result achieves a rate of $\mathcal{O}\left(1/(mn)\right)$, leveraging all $mn$ samples and matching the rate of GlobalTrain on the IID data. As the client number could be extremely large ($10^{5}$ devices in chen2023fs), order of $m$ is non-trivial. During the transition period when $m^{-1}n^{-1/2}<R<n^{-1/2}$, we achieve a rate of $\mathcal{O}(R^{2})$, again matching the lower bound, and is strictly faster than the rate $\mathcal{O}(n^{-1/2}R)$ established in Chen et al. (2023b).